# Faithful global convergence for the rescaled Consensus–Based Optimization

### v1

@@ 3.2. Multiple–minimizer
It can be shown in fact that the claim (3.9) holds with an equality instead of an inequality. The proof of the other inequality relies on a selection principle; see [kouhkouh2024viscous, Proposition 3.1].
Corollary 3.12.
Let $Y_{t}^{\alpha}$ be a random variable whose probability distribution is $\eta_{t}^{\alpha}$ defined by
(3.14) $\eta_{t}^{\alpha}(A):=\frac{\int_{A}e^{-\alpha f(x)}\rho_{t}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{t}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}),\quad\rho_{t}=\mathrm{Law}({\overline{X}}_{t}).$
Then in the situation of Theorem 3.9, we have
$\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0.$
In particular, we have
$\lim\limits_{\alpha\to\infty}\lim\limits_{t\to\infty}\,\mathbb{P}\big{(}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})=0\big{)}=1.$
Note that the latter convergence holds independently of the initial distribution of $Y_{0}^{\alpha}\sim\eta_{0}^{\alpha}$ defined with $\rho_{0}=\mathrm{Law}({\overline{X}}_{0})\in\mathscr{P}_{4}({\mathbb{R}}^{d})$.
Proof.
Let $\delta>0$ be fixed. Markov’s inequality yields
$\mathbb{P}\big{(}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})>\delta\big{)}\leq\frac{\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}}{\delta}.$
Then it suffices to show that $\lim\limits_{\alpha\to\infty}\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0$.
Using the definition of $Y_{t}^{\alpha}$, we have
$\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}$ $=\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
$\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
where the inequality follows from (3.9).
Letting $\alpha$ be fixed, we introduce the notation:
$F_{t}:L^{\infty}({\mathbb{R}}^{d},\mathcal{M})\to\mathbb{R},\quad\quad F_{t}(\Delta):=\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{t}^{\alpha}(\mbox{\rm d}y)$
and, analogously, we define
$F_{*}:L^{\infty}({\mathbb{R}}^{d},\mathcal{M})\to\mathbb{R},\quad\quad F_{*}(\Delta):=\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We shall first prove that the sequence $\{F_{t}\}_{t>0}$ converges uniformly to $F_{*}$ when $t\to\infty$. Consequently we can exchange the minimization and the limit $t\to\infty$, and conclude as in the proof of Corollary 3.10.
Step 1. (Uniform convergence)
Recalling the definition of $\eta_{t}^{\alpha}$ in (3.14), we can express $F_{t}$ such that
$F_{t}(\Delta)=\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)$
where $h_{\Delta}(y)=|y-\Delta(y)|e^{-\alpha f(y)}\quad\text{ and }\quad g(\rho^{\alpha}_{t})=\int_{{\mathbb{R}}^{d}}e^{-\alpha f(x)}\,\rho_{t}^{\alpha}(\mbox{\rm d}x).$
Analogously, we can also write
$F_{*}(\Delta)=\frac{1}{g(\rho_{*}^{\alpha})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y).$
Thus we have
$|F_{t}(\Delta)-F_{*}(\Delta)|$ $\leq\left|\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)-\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\quad\quad\quad+\left|\frac{1}{g(\rho^{\alpha}_{t})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)-\frac{1}{g(\rho_{*}^{\alpha})}\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\leq\frac{1}{g(\rho^{\alpha}_{t})}\left|\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho^{\alpha}_{t}(\mbox{\rm d}y)-\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y)\right|$
$\quad\quad\quad+\left|\frac{1}{g(\rho^{\alpha}_{t})}-\frac{1}{g(\rho_{*}^{\alpha})}\right|\int_{{\mathbb{R}}^{d}}h_{\Delta}(y)\,\rho_{*}^{\alpha}(\mbox{\rm d}y).$
The set $\mathcal{M}$ being compact, one can find a constant $C>0$ such that for all $\Delta\in L^{\infty}(\mathbb{R}^{d},\mathcal{M})$ we have $|h_{\Delta}(y)|\leq C(1+|y|)$. Hence, using (2.8) (see also [zhang2023existence, Lemma 2.7]), we can send $t\to\infty$ in each term of the latter inequality and obtain the desired uniform convergence.
Step 2. (Exchange of minimization and limit)
Consider a quasi-minimizer of $F_{*}$, that is, for any $\varepsilon>0$, there exists $\Delta_{\varepsilon}$ such that
(3.15) $F_{*}(\Delta_{\varepsilon})\leq\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
The uniform convergence guarantees that for any $\varepsilon>0$, there exists $T_{\varepsilon}>0$ such that for all $t\geq T_{\varepsilon}$ it holds
(3.16) $|F_{t}(\Delta)-F_{*}(\Delta)|\leq\varepsilon\quad\forall\,\Delta.$
In particular, when the latter is evaluated in $\Delta_{\varepsilon}$, one has
$F_{t}(\Delta_{\varepsilon})\leq F_{*}(\Delta_{\varepsilon})+\varepsilon$
which together with (3.15) yields
$F_{t}(\Delta_{\varepsilon})\leq 2\,\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
But $\min\limits_{\Delta}F_{t}(\Delta)\leq F_{t}(\Delta_{\varepsilon})$. Therefore we have
$\min\limits_{\Delta}F_{t}(\Delta)\leq 2\,\varepsilon+\min\limits_{\Delta}F_{*}(\Delta).$
Taking first the limit $t\to\infty$, then letting $\varepsilon\to 0$, yield $\lim\limits_{t\to\infty}\min\limits_{\Delta}F_{t}(\Delta)\leq\min\limits_{\Delta}F_{*}(\Delta)$.
We now prove the converse inequality. Given $t>0$, let us consider $\{\Delta_{n,t}\}_{n\in\mathbb{N}}$ a minimizing sequence for $F_{t}(\cdot)$, that is
(3.17) $\lim\limits_{n\to\infty}F_{t}(\Delta_{n,t})=\min\limits_{\Delta}F_{t}(\Delta).$
The inequality (3.16) evaluated in $\Delta_{n,t}$ yields (when choosing $t\geq T_{\varepsilon}$)
$F_{*}(\Delta_{n,t})\leq\varepsilon+F_{t}(\Delta_{n,t}).$
But $\min\limits_{\Delta}F_{*}(\Delta)\leq F_{*}(\Delta_{n,t})$. Therefore we have
$\min\limits_{\Delta}F_{*}(\Delta)\leq\varepsilon+F_{t}(\Delta_{n,t}).$
We can now first take the limit $n\to\infty$ using (3.17), then we let $t\to\infty$, which yield $\min\limits_{\Delta}F_{*}(\Delta)\leq\varepsilon+\lim\limits_{t\to\infty}\min\limits_{\Delta}F_{t}(\Delta)$. The conclusion follows by sending $\varepsilon\to 0$.
Step 3. (Conclusion)
We have shown that
$\lim\limits_{t\to\infty}\,\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}F_{t}(\Delta)=\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}F_{*}(\Delta)$
which then yields
$\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}$ $\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$
$\leq\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\quad\quad\forall\,\Delta\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M}).$
In particular we can evaluate the latter inequality in $\underline{\Delta}$ defined in (3.11), and conclude as in Corollary 3.10 that
$\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{t}^{\alpha},\mathcal{M})\big{]}=0.$
@@ 4. Conclusion
under suitable conditions on the parameters, for example
$\delta=1,\quad\kappa<\frac{1}{2}(1+b_{2})^{-1},\quad\text{ and }\quad\lambda>16\,\sigma^{2}$
@@
where $b_{2}$ is a constant depending on the function $f(\cdot)$ to be minimized (see Assumption 2.1 and Lemma 2.2). We do not make any assumptions on the initial distribution ${\overline{X}}_{0}\sim\rho_{0}$ besides satisfying $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$.
Recalling (2.8) obtained from [zhang2023existence, Lemma 2.7], for any $|\varphi(x)|\leq C(1+|x|)$ it holds
$\lim_{t\to\infty}\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{t}^{\alpha}(x)=\int_{{\mathbb{R}}^{d}}\varphi(x)\,\mbox{\rm d}\rho_{*}^{\alpha}(x)\,$
where $\rho_{*}^{\alpha}$ is the unique invariant probability measure of the rescaled CBO. When combined with (3.4), we obtain
@@
(4.1) $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}\int_{{\mathbb{R}}^{d}}x\,\mbox{\rm d}\rho_{t}^{\alpha}(x)\,=\,x_{*}$
where $x_{*}$ is the unique global minimizer of $f(\cdot)$.
When $f(\cdot)$ has multiple minimizers, we obtain the concentration of the weighted probability measures
$\eta_{*}^{\alpha}(A)=\frac{\int_{A}e^{-\alpha f(x)}\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{*}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}).$
on the set of global minimizers, in the limit $\alpha\to\infty$. Moreover, if we let $Y_{*}^{\alpha}$ be a random variable with probability distribution $\eta_{*}^{\alpha}$, and $Y_{t}^{\alpha}$ be a random variable whose probability distribution is
$\eta_{t}^{\alpha}(A)=\frac{\int_{A}e^{-\alpha f(x)}\rho_{t}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{t}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d}),\quad\rho_{t}=\mathrm{Law}\left({\overline{X}}_{t}\right),$
then we have shown that it converges in mean, hence also in probability, to the set of global minimizers $\mathcal{M}$. More precisely we have

### v2

@@ 4.2. Multiple–minimizer
A result more general than (4.10) can be found in [41, Proposition 3.1].
To validate Corollary 4.10, we apply our CBO dynamics to the two-dimensional objective function
$f(x,y):=\big((x-1)^{2}+(y-1)^{2}\big)\big((x+1)^{2}+(y+1)^{2}\big),$
which possesses two global minimizers at $(1,1)$ and (1,1)(-1,-1). In Figure 2, we plot the trajectory of the consensus points $\mathfrak{m}_{\alpha}(\rho^{N}_{t})$ generated by the CBO algorithm. The results show that, depending on the initial particle distribution, the method converges to one of the two global minimizers, thereby confirming Corollary 4.10. The simulation parameters are chosen as $N=1000$, $T=4$, $\mathrm{d}t=0.01$, $\lambda=1$, $\sigma=0.3$, $\kappa=0.9$, and $\delta=0.8$.
Figure 2. Numerical test on the objective function $f(x,y)$, which has two global minimizers at $(1,1)$ and (1,1)(-1,-1).
Left: The particles are initialized uniformly in $2,-1]\times[3,4]$, a region distant from the global minimizers. The CBO method successfully converges to the global minimizer $(1,1)$.
Right: The particles are initialized uniformly in $3,-2]\times[1,2]$. The CBO method successfully converges to the global minimizer (1,1)(-1,-1).
5. Numerical aspects
In this section, we investigate numerical aspects of the rescaled CBO, with particular emphasis on the sensitivity of the algorithm to the parameters $\kappa$ and $\delta$.
Our results indicate that the rescaled CBO offers significant advantages over the standard CBO, especially when the initial particle distribution does not contain the global minimizer.
To this end, we first apply the rescaled CBO dynamics using the standard Euler–Maruyama scheme to the two-dimensional Rastrigin function $R(\cdot,\cdot)$, which features a unique global minimizer at $x_{*}=(1,1)^{\top}$
$R(x_{1},x_{2})=20+\left(x_{1}^{2}-10\cos(2\pi x_{1})\right)+\left(x_{2}^{2}-10\cos(2\pi x_{2})\right).$
The particles are initially uniformly distributed in the square $2,3]\times[2,3]$, situated away from $x_{*}$ to test the algorithm’s global search capabilities. We fix the parameters at $\lambda=1,\sigma=0.5,\alpha=10^{15},T=100,$ and $\Delta t=0.01$ to evaluate the sensitivity of the success rate across various choices of $\kappa$ and $\delta$. In Table 3, the number of particles is $N=100$. In Table 3, it is $N=500$. A test is deemed successful if $|\mathfrak{m}_{\alpha}(\rho_{T}^{N})-x_{*}|<0.05$; otherwise, it fails. The success rate is calculated based on $100$ test runs.
$\delta$ $\kappa$ Rate 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
0 0 0 0 0 0 0 0 0 0 0
0.5 0 0 0 0 0 0 0.03 0.98 1.00 1.00
1.0 0.02 0.03 0.05 0.08 0.08 0.19 0.49 0.87 0.96 0.98
1.5 0.05 0.06 0.13 0.16 0.31 0.38 0.52 0.72 0.84 0.79
2.0 0.19 0.18 0.21 0.26 0.30 0.33 0.39 0.61 0.68 0.60
2.5 0.16 0.21 0.15 0.29 0.27 0.40 0.43 0.54 0.51 0.49
3.0 0.14 0.17 0.17 0.15 0.25 0.22 0.37 0.44 0.42 0.42
3.5 0.21 0.26 0.20 0.27 0.30 0.37 0.25 0.25 0.35 0.33
4.0 0.15 0.19 0.26 0.20 0.24 0.22 0.28 0.36 0.32 0.32
4.5 0.16 0.15 0.18 0.16 0.31 0.24 0.27 0.23 0.34 0.22
5.0 0.11 0.21 0.20 0.18 0.21 0.14 0.28 0.26 0.31 0.33
Table 2. Success rate for $N=100$.
$\delta$ $\kappa$ Rate 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
0 0 0 0 0 0 0 0 0 0 0
0.5 0 0.01 0 0.03 0.02 0.04 1.00 1.00 1.00 1.00
1.0 0.08 0.13 0.15 0.31 0.61 0.87 0.97 1.00 1.00 1.00
1.5 0.30 0.29 0.40 0.56 0.76 0.91 0.94 0.99 0.99 1.00
2.0 0.46 0.37 0.50 0.60 0.72 0.81 0.92 0.93 0.99 0.98
2.5 0.40 0.37 0.50 0.61 0.66 0.82 0.84 0.90 0.94 0.91
3.0 0.47 0.39 0.51 0.63 0.67 0.75 0.74 0.83 0.81 0.81
3.5 0.49 0.48 0.51 0.44 0.53 0.64 0.67 0.67 0.76 0.72
4.0 0.40 0.43 0.49 0.59 0.48 0.62 0.65 0.57 0.72 0.71
4.5 0.40 0.43 0.40 0.52 0.46 0.50 0.58 0.53 0.55 0.62
5.0 0.34 0.51 0.47 0.37 0.48 0.48 0.54 0.43 0.55 0.51
Table 3. Success rate for $N=500$.
As observed in the table, optimal performance is generally achieved by choosing a large $\kappa$ and a small $\delta$ (the upper-right region of the table). Notably, the standard CBO configuration ($\kappa=1,\delta=0$) fails completely with a success rate of 0. These results further validate the motivation for introducing the scaling parameter $\kappa\in(0,1)$ and the non-degenerate diffusion $\delta>0$ to enhance convergence. Furthermore, one can observe in Table 3 that increasing the number of particles generally improves the success rate.
We further apply the CBO dynamics to the Ackley function
$A(x):=-20\exp\!\left(-0.2|x-x_{*}|\right)-\exp\!\left(\frac{1}{d}\sum_{i=1}^{d}\cos(2\pi(x_{i}-x_{*,i}))\right)+e+20,$
in dimension $d=20$, where the unique global minimizer is $x_{*}=(1,\dots,1)^{\top}$.
Figure 3. We test the Ackley function in dimension $d=20$.
Left: The particles are initialized uniformly in $2,3]^{20}$ (which excludes $x_{*}$). The rescaled CBO ($\kappa=0.9$, $\delta=0.8$) converges to the global minimizer, while the standard CBO stalls at a local minimum.
Right: The particles are drawn from $\mathcal{N}((2,\dots,2)^{\top},\mathds{I}_{20})$ (whose support includes $x_{*}$). The rescaled CBO (with $\kappa=0.9$ and $\delta=0.8$) still outperforms the standard CBO.
In Figure 3, we fix the parameters $\lambda=1$, $\sigma=0.3$, $\alpha=10^{15}$, N=500  [^0] N=500\,000, $T=20$, and $\Delta t=0.01$. In Figure 3-(Left), the particles are initially sampled uniformly from the hypercube $2,3]^{20}$, which does not contain $x_{*}$. The plot shows the evolution of the error $\log(|\mathfrak{m}_{\alpha}(\rho_{t}^{N})-x_{*}|)$ over time $t$. One observes that the standard CBO becomes trapped in a local minimizer, whereas the rescaled CBO with $\kappa=0.9$ and $\delta=1$ successfully converges to the global minimizer. In Figure 3-(Right), we consider an initial distribution $\mathcal{N}((2,\dots,2)^{\top},\mathds{I}_{20})$, whose support includes $x_{*}$. Even in this more favorable setting, the rescaled CBO demonstrates better performance, highlighting its robustness and effectiveness beyond merely escaping poor initialization.
Several additional numerical experiments (using dimensions up to 20 and a variety of objective functions) are reported in [31, Section 5]. In fact, the optimization problem studied in [31] is multi-level such as
$\min\limits_{x\in\mathbb{R}^{n}}\,F(x,y)$
$\quad\text{s.t.}\quad y\in\operatorname*{argmin}_{y\in\mathbb{R}^{m}}\,G(x,y),$
where $F,G:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ for some $n,m\in\mathbb{N}$. In order to fit it into the present framework, it suffices to take the objective function $F$ to be constant (a simplification) and drop the dependence on its corresponding $x$ variable (by setting it to a fixed constant) in the objective function $G$. The resulting optimization problem becomes
$\left.\begin{aligned} &\min\limits_{x\in\mathbb{R}^{n}}\quad 1\\ &\quad\text{s.t.}\quad y\in\operatorname*{argmin}_{y\in\mathbb{R}^{m}}\,G(0,y)\end{aligned}\quad\right\}\quad\Longleftrightarrow\qquad\min\limits_{y\in\mathbb{R}^{m}}\,\overline{G}(y)$
where $\overline{G}(y):=G(0,y)$. Hence, the multi-scale dynamics [31, equation (3.1)] reduces to $(0,\mathds{Y}_{t})$ where the process $\mathds{Y}$ solves (1.6). With this simplification, the model in [31] encompasses the one considered here, and the remainder of [31, Section 5] is fully compatible with our setting. Consequently, the numerical tests and the results obtained therein apply directly and would be identical in our case. See also [31, Remark 5.1].
@@ 6.1. Summary of main findings
under suitable conditions on the parameters, for example those in Remark 3.6. We do not make any assumptions on the initial distribution ${\overline{X}}_{0}\sim\rho_{0}$ besides satisfying $\rho_{0}\in\mathscr{P}_{4}({\mathbb{R}}^{d})$. Supposing that $f(\cdot)$ satisfies Assumption 2.1:
  • We proved that such a process admits a unique invariant probability measure $\rho^{\alpha}_{*}$, and its law enjoys a $\mathbb{W}_{2}$–exponential contraction (3.1).
  • When $f(\cdot)$ has a unique minimizer $x_{*}$, we proved that
    $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})=\lim_{\alpha\to\infty}\;\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})=x_{*},$
@@
    and $\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\;\kappa^{-1}\int_{{\mathbb{R}}^{d}}x\,\mbox{\rm d}\rho^{\alpha}_{*}(x)\,=\,x_{*},$
  • When $f(\cdot)$ has multiple minimizers as $\mathcal{M}$, we proved that
    $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}),\,\mathcal{M})=\lim\limits_{\alpha\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})=0,$